how many gallons of 70% antifreeze needs to be added to 10 gallons of 30% antifreeze to get a 45% mix?
Answer by stanbon(75887)

how many gallons of 70% antifreeze needs to be added to 10 gallons of 30% antifreeze to get a 45% mix?
---
Equation:
active + active = active
0.70x + 0.30*10 = 0.45(x+10)
Multiply thru by 100 to get:
70x + 30*10 = 45x + 45*10
25x = 15*10
x = 6 gallons (amt. of 70% solution needed)
==================
Cheers,
Stan H.
=============